Practice Problem 2: Crash Course in Optimization: Part II 1. Consider a monopolist who pays an output tax $t for each unit ofhave, it's That's really all I have, in fact, I've figured out the questions A,B,C, but I don't know how to approach the rest of the question.
Answer #1

D).

Consider the given problem here the market demand curve is “P=a-b*q”, => the TR = P*q = a*q-b*q^2.

=> TR = a*q-b*q^2, => MR = d(TR)/dq = a-2b*q, = > MR = a-2b*q.

Now, the cost function is “C=c*q^2”, => C’() = 2c*q > 0” for all “q > 0” and “C’’() = c > 0”.

Now, the profit function is given below.

=> A = TR - TC - t*q, => the FOC for maximization is “dA/dq = 0”.

=> dA/dq = d(TR)/dq – d(TC)/dq - t = 0, => MR – MC - t = 0, => MR=MC+t.

=> a-2b*q=2c*q+t, => a-t = 2(c+b)*q, => q = (a-t)/2(b+c), the equilibrium quantity.

The above quantity expression can also be written as follows.

=> q = a/2(b+c) - t/2(b+c), => dq/dt = (-1)/(b+c) < 0, => as tax increases implied the equilibrium quantity decreases and vice versa.

E).

The market equilibrium price is given by, => P = a-b*q, => P = a-b*[(a-t)/2(b+c)]

=> P = [2a*(b+c) - b*(a-t)]/2(b+c), => P = [2ab + 2ac - a*b + b*t]/2(b+c).

=> P = [ab + 2ac + bt]/2(b+c), => dP/dt = b/2(b+c) > 0, => as the tax increases implied the equilibrium price increases and vice versa.

Now, as we know that “b < b+c”, => b/2(b+c) < 1, => as tax increases by “1 rupees” the equilibrium price increases by less than 1 rupees.

F).

Now, let’s assume that the cost function is linear, => the equilibrium rice is “P = (a+c+t)2”, => dP/dt = 1/2. So, as the tax increases by “1 rupees”, => the equilibrium price increases by “1/2 rupees”.

Now, if the cost function is convex, => the equilibrium rice is “P = [ab + 2ac + bt]/2(b+c), => dP/dt = b/2(b+c) = (1/2)*(b/b+c) < 1/2. So, as the tax increases by “1 rupees”, => the equilibrium price increases by less than “1/2 rupees”.

So, the price increases by more under the linear cost function compare to the convex cost function.